### v2

# Integrating large language models and active inference to understand eye movements in reading and dyslexia

Abstract
We present a novel computational model employing hierarchical active inference to simulate reading and eye movements. The model characterizes linguistic processing as inference over a hierarchical generative model, facilitating predictions and inferences at various levels of granularity, from syllables to sentences. Our approach combines the strengths of large language models for realistic textual predictions and active inference for guiding eye movements to informative textual information, enabling the testing of predictions. The model exhibits proficiency in reading both known and unknown words and sentences, adhering to the distinction between lexical and nonlexical routes in dual route theories of reading. Notably, our model permits the exploration of maladaptive inference effects on eye movements during reading, such as in dyslexia. To simulate this condition, we attenuate the contribution of priors during the reading process, leading to incorrect inferences and a more fragmented reading style, characterized by a greater number of shorter saccades. This alignment with empirical findings regarding eye movements in dyslexic individuals highlights the model’s potential to aid in understanding the cognitive processes underlying reading and eye movements, as well as how reading deficits associated with dyslexia may emerge from maladaptive predictive processing. In summary, our model represents a significant advancement in comprehending the intricate cognitive processes involved in reading and eye movements, with potential implications for understanding and addressing dyslexia through the simulation of maladaptive inference. It may offer valuable insights into this condition and contribute to the development of more effective interventions for treatment.
keywords:
Hierarchical active inference, Predictive coding, Large Language Models, Reading, Dyslexia
[…]
[CNR]organization=Institute of Cognitive Sciences and Technologies - National Research Council
,addressline=Via San Martino della Battaglia 44, city=Rome, postcode=00185, state=Italy, country=Italy
Introduction
Processing natural language – encompassing understanding, reading, and producing linguistic content – represents a fundamental ability of our species. Extensive research in psychology, neuroscience, linguistics, and machine learning has explored the intricate ways we process natural language.
Neuroscientific studies have revealed that natural language processing is inherently hierarchical, involving multiple brain regions and the integration of various sensory inputs (Capek et al., 2004; Butterfuss and Kendeou, 2018; Price, 2012). This hierarchical processing spans from individual letters, phonemes, and words to complete sentence comprehension, with a hierarchy of brain areas actively maintaining these elements in working memory across multiple time scales (Hasson et al., 2015).
Recent studies have increasingly supported the idea of hierarchical predictive coding as a formal theory describing perception as an inferential process, involving reciprocal exchanges between predictions and prediction errors across brain hierarchies (Rao and Ballard, 1999; Friston, 2005). In particular, various computational neuroimaging studies employing large language models (LLMs) during linguistic tasks have provided compelling evidence for both the predictive nature of language processing and the prediction hierarchies proposed by predictive coding (Goldstein et al., 2022; Schmitt et al., 2021; Baldassano et al., 2017; Zadbood et al., 2017; Chang et al., 2022; Caucheteux et al., 2023; Heilbron et al., 2022).
A recent computational neuroimaging study investigated human electrocorticographic (ECoG) responses to narratives using LLMs trained to predict the next word. Remarkably, the study revealed that like the LLMs, the brain engages in next-word prediction before word onset, computes prediction error signals, and utilizes latent representations of words (embeddings) contextualized based on the sequence of prior words (Goldstein et al., 2022).
Converging evidence emerges from three recent fMRI studies utilizing deep learning models during linguistic tasks. These studies not only confirm the predictive nature of language processing but also lend support to the prediction hierarchies proposed by predictive coding. The first fMRI study (Schmitt et al., 2021) trareferined a LLM to predict the next word at multiple time scales, identifying event boundaries as high surprise (also explored in (Baldassano et al., 2017; Zadbood et al., 2017; Chang et al., 2022)). Analyzing human functional magnetic resonance data during story listening, the study revealed an event-based hierarchy of surprise signals evolving along temporoparietal regions, with surprise signals gating bottom-up and top-down connectivity across neighboring time scales.
Furthermore, another fMRI study provided compelling evidence of the brain’s ability to predict a hierarchy of representations spanning multiple timescales in the future (Caucheteux et al., 2023). Enhancing LLMs with the capability to predict beyond the next word increased their fit with human data. The study also highlighted the hierarchical organization of brain predictions, ranging from temporal cortices predicting shorter-range representations (e.g., the next word) to frontoparietal cortices predicting higher-level, longer-range, and more contextual representations.
Notably, a separate fMRI study reported that evoked brain responses to words are influenced by linguistic predictions and a metric of unexpectedness, closely aligning with the hierarchical scheme of active inference, where lower-level predictions are informed by higher-level predictions (Heilbron et al., 2022). Additionally, the study demonstrated that these hierarchical predictions can be well-aligned with standard levels of analysis in psycholinguistics, including meaning, grammar, words, and speech sounds, reinforcing the validity of the standard decomposition.
Taken together, these studies, alongside others (Kutas and Hillyard, 1984; Weissbart et al., 2020; Koskinen et al., 2020; Willems et al., 2016; Frank et al., 2015), provide compelling support for the significance of prediction and hierarchical predictive coding in language processing. Moreover, they underscore the growing relevance of LLMs in comprehending linguistic processing (Goldstein et al., 2022).
Despite these advancements, the above studies have primarily focused on LLMs that passively receive sensory information, rather than actively searching for it. However, linguistic tasks, such as reading written text and listening to speech, are inherently active processes (Ferro et al., 2010; Norris, 2006; Friston et al., 2021). For example, during reading, eye movements (saccades) actively guide attention to relevant parts of the text, rather than processing every piece of text linearly. This active reading process suggests that saccades play a crucial role in hypothesis testing, selecting informative parts of the text to test predictions (Ferro et al., 2010; Norris, 2006; Donnarumma et al., 2017; Friston et al., 2012).
Moreover, eye movement studies during reading have proven to be valuable in understanding reading difficulties, such as dyslexia. Dyslexia, a common reading disorder affecting 5-10% of the population, is associated not only with atypical brain activation patterns during language processing (Price, 2012) but also with atypical eye movement patterns, such as increased numbers of forward saccades and decreased saccade lengths compared to control groups (Zoccolotti et al., 2005; De Luca et al., 1999; Lyon et al., 2003; Franzen et al., 2021).
The connection between prediction-based language processing and the perceptual challenges observed in dyslexia has been explored in previous studies like (Jaffe-Dax et al., 2015). These investigations suggest that dyslexia could be characterized as a computational difficulty in effectively combining prior information, such as implicit memory of previous words, with noisy observations like the currently perceived word. Essentially, when prior information is given excessively low weighting relative to its internal noise, the time required to form a coherent understanding of text (or to perform auditory discrimination, as in (Jaffe-Dax et al., 2015)) significantly increases because novel observations consistently surprise dyslexics. Through computational simulations, the authors demonstrated that this pattern of results could account for dyslexics’ perceptual deficit. However, the proposed model lacks the capability to actively select information, such as generating eye movements.
Several reading models incorporating eye movements, such as E-Z Reader (Reichle et al., 1998), SWIFT (Engbert et al., 2005; Seelig et al., 2020; Rabe et al., 2021), Über-Reader (Reichle, 2021; Veldre et al., 2020), Glenmore (Reilly and Radach, 2002, 2006), and OB1-Reader (Snell et al., 2018) significantly advanced our understanding of the role of reading dynamics. These models have successfully replicated numerous empirical findings, highlighting how word-level attributes like length, frequency, and predictability impact reading dynamics. However, these models focus largely on data regression and they do not provide a mechanistic explanation of the cognitive processes integral to reading, such as for example how saccades are generated.
In this paper, we propose a novel computational model that unifies hierarchical predictive processing and hypothesis testing during reading by integrating the LLM BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019) with active inference (Friston, 2010; Parr et al., 2022). Our model views reading as an active (Bayesian) inference problem, employing a hierarchical generative model to represent causal relationships between textual elements at different levels (letters, syllables, words, and sentences). By generating predictions at each level and testing them through saccades, our model actively simulates reading.
The model incorporates three significant insights discussed earlier. Firstly, it conceptualizes linguistic processing as inference, employing a hierarchical generative model that allows for predicting and inferring at different timescales, such as syllables, words, and sentences. By integrating the LLM BERT (Devlin et al., 2019) at the highest hierarchical level, our model can handle realistic reading tasks effectively, processing sentences of arbitrary length. Additionally, thanks to its hierarchical structure, our model can read text both word-by-word, a common approach in LLMs (i.e., the lexical route), and syllable-by-syllable, akin to the nonlexical route in dual route theories of reading (Coltheart, 2005).
Secondly, the model utilizes its generative capabilities not only for language recognition and prediction, similar to LLMs, but also for simulating eye movements and saccades. In our model, reading involves an active, hypothesis testing process: the model generates saccades to the most informative parts of the text to validate its predictions and disambiguate among competing hypotheses about the content being read (Ferro et al., 2010; Baldassano et al., 2017; Donnarumma et al., 2017; Friston et al., 2012; Norris, 2006).
Thirdly, by reducing the influence of prior information in the inference (Jaffe-Dax et al., 2015), our model produces simulated eye movements during reading that align with empirical findings in dyslexia (Zoccolotti et al., 2005; De Luca et al., 1999; Hutzler and Wimmer, 2004; Rayner, 1983). This aspect allows the model to generate eye movement patterns consistent with dyslexic individuals’ reading impairments.
This model allows us to gain deeper insights into the integration of information during reading, the role of saccades in testing predictions, and the potential impact of maladaptive predictive processing in reading deficits associated with dyslexia.
Furthermore, its hierarchical composition from primitive content unveils phenomena challenging to investigate with other models, such as reading unfamiliar or unknown words.
In the following sections, we demonstrate the model’s capabilities through simulations of word reading (Simulation 1), sentence reading (Simulation 2), reading unknown words and sentences (Simulation 3), and reading with prior information about the topic (Simulation 4).
1 Methods
In this section, we provide a summary description of the model, a brief introduction to active inference (Parr et al., 2022), and a detailed discussion of the hierarchical active inference reading model used in this paper.
1.1 Brief explanation of the model
We present a novel hierarchical active inference model for reading and eye movements during reading (Friston, 2010). Consistent with evidence of hierarchical language processing in the brain (Heilbron et al., 2022), our model consists of three levels representing syllables, words, and sentences, as illustrated in Figure 1(a).
At each level, the inference process involves integrating three types of messages: bottom-up messages from nodes at the level below (conveying observations), top-down messages from nodes at the level above (conveying predictions), and lateral messages from nodes representing the previous timestep at the same level (providing memory). For instance, the inference about the current syllable is informed by bottom-up observations about the currently observed letter, top-down predictions from the currently inferred word, and lateral information about the previous syllable. Similarly, the inference about the current word relies on bottom-up observations (i.e., the currently estimated word), top-down predictions from the level above (i.e., at the sentence level), and information about the previous word.
Figure 1: The hierarchical active inference model of reading and eye movements. (a) The hierarchical model comprises three layers for syllables (Level 1), words (Level 2), and sentences (Level 3), represented using the formalism of Partially Observable Markov Decision Processes (POMDPs). Empty nodes indicate hidden variables, filled nodes represent observations, and edges depict probabilistic relations between the nodes, as detailed in B. (b) Model parameters. (c) The sequence of saccades generated by the model while reading an example sentence. Yellow dots indicate the positions of the saccades in the text, and colored lines trace the sequence of saccades from red to yellow. (d) Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time. The blue circles at the bottom level indicate saccades. Once the first level confidently infers a syllable (or when the maximum level of iterations is reached), it sends a bottom-up signal to the second level (indicated by the orange vertical bar) to aid in word inference. Similarly, the second level sends a signal to the third level (indicated by the yellow vertical bar) upon confidently inferring a word. Please refer to the main text for further explanation.
To ensure that our model is capable of reading realistic text, we utilize the large language model BERT (Devlin et al., 2019) to generate a probability distribution over the 100 most likely sentences consisting of 4-8 words. Importantly, this process can be iterated to read sentences of varying lengths, providing the flexibility to handle texts of different complexities.
Lastly, our model has the capability to determine where to make (forward or backward) saccades and select the next letter observation (Donnarumma et al., 2017). All three levels of the model are involved in the process of selecting where to direct the next saccade, by selecting the word (Level 3), syllable (Level 2), and letter (Level 1) to attend to. The selection is guided by considering the information gain of lexical elements at all levels, see Section 1 for further details.
Figure 1(b) visually represents the saccades generated by our model while reading an example sentence: Active inference model of eye movements. The model demonstrates uncertainty about the first word (active) and therefore reads it using two saccades to the first (a) and fourth (i) letters. It follows the same process for the subsequent two words (inference and model). However, at this stage, it skips the next two words, as it perceives the most informative word to be the one following (movements). This mechanism showcases the model’s ability to dynamically select the most relevant information during the reading process.
Figure 1(c) illustrates the probabilities of correctly recognized syllables (blue), words (red), and sentences (yellow) during the reading process. Over time, these probabilities approach one, indicating successful recognition. Additionally, the figure demonstrates that the levels operate at different time scales, with lower levels processing information more quickly than higher levels. This temporal separation arises because each level sends messages to the level above only after accumulating sufficient confidence (as shown by the vertical bars in the figure), a process that typically requires multiple rounds of inference (Friston et al., 2017).
1.2 Brief introduction to Active Inference
Active Inference is a framework that models an agent’s action-perception loop by minimizing variational free energy (Parr et al., 2022). The key objective is to minimize free energy, and to achieve this, the agent possesses a generative model (like the one illustrated in Figure 1), which captures the joint probability of the stochastic variables (hidden states and observations) using the formalism of probabilistic graphical models (Bishop and Nasrabadi, 2006). The agent’s perception involves making inferences about the hidden states based on the observed sensory inputs, while action involves selecting actions that can change the hidden states and optimize the model’s predictions. The generative model for active inference is defined as follows:
$P\left(o_{0:K},s_{0:K},u_{1:K},\gamma|\boldsymbol{\Theta}\right)=$
$P\left(\gamma|\boldsymbol{\Theta}\right)P\left(\pi|\gamma,\boldsymbol{\Theta}\right)P\left(s_{0}|\boldsymbol{\Theta}\right)\prod_{t=0}^{K}P\left(o_{t}|s_{t},\boldsymbol{\Theta}\right)P\left(s_{t+1}|s_{t},\pi_{t},\boldsymbol{\Theta}\right)$ (1)
where
  • 1.
    $P\left(o_{t}|s_{t},\boldsymbol{\Theta}\right)=\mathbf{A}$,
  • 2.
    ${P\left(s_{t+1}|s_{t},\pi_{t},\boldsymbol{\Theta}\right)=\mathbf{B}\left(u_{t}=\pi_{t}\right)}$,
  • 3.
    ${P\left(\pi_{t}|\gamma,\boldsymbol{\Theta}\right)=\sigma\left(\ln\mathbf{E}-\gamma\cdot\mathbf{G}^{\boldsymbol{\Theta}}\left(\pi_{t}\right)\right)}$,
  • 4.
    $P\left(\gamma,\boldsymbol{\Theta}\right)\sim\Gamma\left(\alpha,\beta\right)$, and
  • 5.
    $P\left(s_{0}|\boldsymbol{\Theta}\right)=\mathbf{D}$.
The set $\boldsymbol{\Theta}=\{\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D},\mathbf{E},\alpha,\beta\}$ parametrizes the generative model:
  • 1.
    The (likelihood) matrix $\mathbf{A}$ encodes the relations between the observations $\mathrm{O}$ and the hidden causes of observations $\mathrm{S}$.
  • 2.
    The (transition) matrix $\mathbf{B}$ defines how hidden states evolve over time $t$, as a function of a control state (action) $u_{t}$; note that a sequence of control states $u_{1}$,$u_{2}$$\dots$,$u_{t}$$\dots$ defines an action policy (or policy for short) $\pi_{t}$.
  • 3.
    The matrix $\mathbf{C}$ is an a-priori probability distribution over observations (and typically encodes the agent’s preferences).
  • 4.
    The matrix $\mathbf{D}$ is the prior belief about the initial hidden state, before receiving any observation.
  • 5.
    $\mathbf{E}$ encodes a prior over the policies (reflecting habitual components of action selection).
  • 6.
    $\gamma\in\mathbb{R}$ is a precision that regulates action selection and is sampled from a $\Gamma$ distribution, with parameters $\alpha$ and $\beta$.
In active inference, the process of perception involves estimating hidden states based on observations and previous hidden states. At the start of the simulation, the model has access to an initial state estimate $s_{0}$ through $\mathbf{D}$ and receives an observation $o_{0}$ that helps refine the estimate by using the likelihood matrix $\mathbf{D}$. Subsequently, for each time step $t=1,\dots,K$, the model infers its current hidden state $s_{t}$ by considering the transitions determined by the control state $u_{t}$, as specified in $\mathbf{B}$. Active inference employs an approximate posterior over (past, present, and future) hidden states and parameters ($s_{0:K},u_{1:K},\gamma$). This approach utilizes a factorized form of variational inference, which corresponds to a framework developed in physics known as mean field theory (Parisi, 1988). Using a mean field approximation, namely assuming that all variables are independent, the factorized approximated posterior can be expressed as:
$Q\left(s_{0:K},u_{1:K},\gamma\right)=Q\left(\pi\right)Q\left(\gamma\right)\prod_{t=0}^{K}Q\left(s_{t}|\pi_{t}\right)$ (2)
where the sufficient statistics are encoded by the expectations $\boldsymbol{\mu}=\left(\mathbf{\tilde{s}}^{\boldsymbol{\pi}},\boldsymbol{\pi},\boldsymbol{\gamma}\right)$, with $\mathbf{\tilde{s}}^{\boldsymbol{\pi}}=\mathbf{\tilde{s}}_{0}^{\boldsymbol{\pi}},\dots,\mathbf{\tilde{s}}_{K}^{\boldsymbol{\pi}}$. Following a variational approach, the distribution in (2) best approximates the posterior when its sufficient statistics $\mu$ minimise the free energy of the generative model, see (Parr et al., 2022). This condition holds when the sufficient statistics are:
$\boldsymbol{s}_{t}^{\boldsymbol{\pi}}$ $\approx\sigma\left(\ln\mathbf{A}\cdot o_{t}+\ln\left(\mathbf{B}\left(\pi_{t-1}\right)\cdot\boldsymbol{s}_{t-1}^{\boldsymbol{\pi}}\right)\right)$ (3a)
$\boldsymbol{\pi}$ $=\sigma\left(\ln\mathbf{E}-\boldsymbol{\gamma}\cdot\mathbf{G}\left(\pi_{t}\right)\right)$ (3b)
$\boldsymbol{\gamma}$ $=\frac{\alpha}{\beta-\boldsymbol{G}\left(\boldsymbol{\pi}\right)}$ (3c)
Action selection in the active inference framework involves choosing a policy, represented by a sequence of control states $u_{1},u_{2},\dots,u_{t}$, that is expected to minimize free energy most effectively in the future. The policy distribution $\boldsymbol{\pi}$ is defined by the Softmax function $\sigma(\cdot)$ and the expected free energy (EFE) of the policies, denoted by $\mathbf{G}$. The EFE incorporates goal-directed components of action selection. Additionally, the precision term $\gamma$ plays a role in encoding the confidence of beliefs concerning $\mathbf{G}$.
The EFE $\mathbf{G}(\pi_{t})$ of each policy $\pi_{t}$ is defined as:
$\mathbf{G}\left(\pi_{t}\right)=\sum_{\tau=t+1}^{K}D_{KL}\left[Q\left(o_{\tau}|\pi\right)\parallel P\left(o_{\tau}\right)\right]+\mathbb{E}_{\tilde{Q}}\left[H\left[P\left(o_{\tau}|s_{\tau}\right)\right]\right]$ (4)
where $D_{KL}\left[\cdot\parallel\cdot\right]$ and $H\left[\cdot\right]$ are, respectively, the Kullback-Leibler divergence and the Shannon entropy, $Q\left(o_{\tau},s_{\tau}|\pi\right)\triangleq P\left(o_{\tau},s_{\tau}\right)Q\left(s_{\tau}|\pi\right)$ is the predicted posterior distribution, $Q\left(o_{\tau}|\pi\right)=\sum_{s_{\tau}}Q\left(o_{\tau},s_{\tau}|\pi\right)$ is the predicted outcome, $P\left(o_{\tau}\right)$ is a categorical distribution representing the preferred outcome and encoded by $\mathbf{C}$, and $P\left(o_{\tau}|s_{\tau}\right)$ is the likelihood of the generative model encoded by the matrix $\mathbf{A}$.
The EFE in (4) can be used as a quality score for the policies and has two terms:
  • 1.
    Expected Cost. The first term of (4) is the Kullback-Leibler divergence between the (approximate) posterior and prior over the outcomes and it constitutes the pragmatic (or utility-maximizing) component of the quality score. This term favours the policies that entail low risk and minimise the difference between predicted ($Q\left(o_{\tau}|\pi\right)$) and preferred ($P\left(o_{\tau}\right)\equiv\mathbf{C}$) future outcomes.
  • 2.
    Expected Ambiguity. The second term of (4) is the expected entropy under the posterior over hidden states and it represents the epistemic (or uncertainty-minimizing) component of the quality score. This term favours policies that lead to states that maximize information gain and diminish the uncertainty of future outcomes $H\left[P\left(o_{\tau}|s_{\tau}\right)\right]$.
After scoring all the policies using the EFE, action selection is performed by drawing over the action posterior expectations derived from the sufficient statistic $\boldsymbol{\pi}$ computed via (3b). Then, the selected action is executed, the model receives a novel observation and the perception-action cycle starts again. See (Parr et al., 2022) for more details.
1.3 Hierarchical Active Inference model for reading
The active inference model employed in this study is structured hierarchically, consisting of three levels denoted by $i\in{1,2,3}$ in our simulations. At each level, the model encodes hidden variables associated with different aspects of textual content: syllables at Level 1, words at Level 2, and sentences at Level 3. This hierarchical organization allows the model to effectively capture and process information at multiple linguistic scales, enabling the inference and generation of syllables, words, and sentences during the reading process. For a visual representation of the model’s architecture, refer to Figure 1(a-b).
The hidden states of each level are obtained by the tensorial product: $S^{(i)}=S^{(i)}_{1}\otimes S^{(i)}_{2}\otimes S^{(i)}_{3}$ among three factors:
  • 1.
    the Content $S^{(i)}_{1}$, i.e., the textual content proper: Syllable, Word or Sentence,
  • 2.
    the Location $S^{(i)}_{2}$, i.e. the position of a portion of the textual content (Location of the Letter in the syllable, Location of the Syllable in the word, and Location of the Word in the Sentence), thus corresponding to a $S^{(i-1)}_{1}$-type content of the next lower level if $i>1$ or to location of a Letter in a Syllable if $i=1$.
  • 3.
    the Topic $S^{(i)}_{3}$, i.e. the context to which each Content variable is associated. Please note that we only use this factor in Simulation 4.
The Content is the crucial element of the chain: it is the text that is obtained from the recursive concatenation of the elements from the level below. A Syllable is a concatenation of $N^{(1)}$ Letters $c^{(0)}$ (which are not hidden states but observations, belonging to a finite Alphabet $S^{(0)}_{1}$, see below). A Word is a concatenation of $N^{(2)}$ syllables $c^{(1)}$ belonging to the set of Syllables $S^{(1)}_{1}$ and a Sentence is concatenation of $N^{(3)}$ Words $c^{(2)}$ belonging to the set $S^{(2)}_{1}$. In general, the $i$-th content $c^{(i)}\in S^{(i)}_{1}$ is a concatenation of $N^{(i)}$ sub-contents $c^{(i-1)}_{n}\in S^{(i-1)}_{1}$ such that
$c^{(i)}=c^{(i-1)}_{1},\dots c^{(i-1)}_{N^{(i)}}$
The observations $O^{(i)}=O^{(i)}_{1}\otimes O^{(i)}_{2}\otimes O^{(i)}_{3}$ consists of the tensorial product among:
  • 1.
    the observation $O^{(i)}_{1}$ of the content of the level below $S^{(i-1)}_{1}$. Note that only Level 1 of the model receives an actual textual observation: namely, a letter $c^{(0)}$ belonging to the Alphabet $S^{(0)}_{1}$. The other two levels receive as observations the content at the level below: the observations at levels 2 and 3 are the inferred syllables and words, respectively;
  • 2.
    the observation $O^{(i)}_{2}$ corresponding to the location $S^{(i)}_{2}$ of the $S^{(i-1)}_{1}$ content;
  • 3.
    the feedback response $r\in O^{(i)}_{3}$, (“correct” or “wrong”) that reports whether the currently inferred content $C^{(i)}$ has been correctly classified (i.e., assigned to the correct topic) or not. We only use this observation at Level 3 (i.e., $O^{3}_{3}$) of Simulation 4, in order to compare models with or without prior information about the topic of the sentence.
The likelihood mapping $p(O^{(i)}|S^{(i)})$ between hidden states $S^{(i)}$ and observations $O^{(i)}$ is specified through the tensor $\mathbf{A}^{(i)}$, defined as the tensorial product $\mathbf{A}^{(i)}=\mathbf{A}^{(i)}_{1}\otimes\mathbf{A}^{(i)}_{2}\otimes\mathbf{A}^{(i)}_{3}$ where
  • 1.
    $\mathbf{A}^{(i)}_{1}$ is a 4-order tensor mapping the hidden states $S^{(i)}$ (3-order tensors) to the observation $O^{(i)}_{1}$ corresponding to the content $S^{(i-1)}_{1}$ if $i>1$ or to the observation of a letter if $i=1$; In our simulations, this probability is typically an identity matrix: the probability of observing content $c^{(i)}\in S^{(i)}_{1}$ if the reader agent is in the corresponding location $l^{(i)}\in S^{(i)}_{2}$ is set to 1. Note that adding noise to the likelihood function would allow modeling the effects of perceptual deficits or poor familiarity with words.
  • 2.
    $\mathbf{A}^{(i)}_{2}$ is a 4-order tensor mapping the hidden states $S^{(i)}$ to observation $O^{(i)}_{2}$ corresponding to the locations $S^{(i)}_{2}$; Also in this case we assume no noise, so the probability of observing a given location when the reader agent eye point to that location is set to 1.
  • 3.
    $\mathbf{A}^{(i)}_{3}$ is a 4-order tensor mapping the hidden states $S^{(i)}$ to the feedback response in $O^{(i)}_{3}$. In our simulation we adopt a binary probability in our simulation: the likelihood is set to 1 if the content belongs to the correct topic and to 0 otherwise. Please note that we only use this factor in Simulation 4.
The mapping $p(s^{(i)}|s^{(i)}_{t-1},\pi_{t})$ between hidden states given the control state u is specified by the tensor $\mathbf{B}^{(i)}$, defined as the tensorial product $\mathbf{B}^{(i)}=\mathbf{B}^{(i)}_{1}\otimes\mathbf{B}^{(i)}_{2}\otimes\mathbf{B}^{(i)}_{3}$ where
  • 1.
    $\mathbf{B}^{(i)}_{1}$ maps the transition between content states at successive time steps. Note that in our simulations, we use transitions models without noise (for the control model) and with noise added at one or more hierarchical levels (for the dyslexic models). For this, when we formalize the mapping $p(c^{(i)}_{t}|c^{(i)}_{t-1},\pi_{t})=\delta^{(i)}$ we use a modulating factor $\delta^{(i)}$, which we set to 1 in the case without noise and to a number below 1 in the case with noise. Noise implies that the prior knowledge about (or the working memory) of the content variable fades away during reading – requiring more evidence and more saccades to solve the reading tasks.
  • 2.
    $\mathbf{B}^{(i)}_{2}$ maps the transitions between locations. We equip the model with the probability $p(l^{(i)}_{t}|l^{(i)}_{t-1},\pi_{t})$ to jump from one location $l^{(i)}\in S^{(i)}_{2}$ to another, without noise (i.e., the act of jumping to the selected location is executed with no errors). While in principle one can jump to any location, here we restrict the jump possibilities, by setting priors over locations $\mathbf{D}^{(i)}$ and policies $\mathbf{E}^{(i)}$, see below.
  • 3.
    $\mathbf{B}^{(i)}_{3}$ maps the probability $p(tp^{(i)}_{t}|tp^{(i)}_{t-1},\pi_{t})$ to jump from one topic $tp^{(i)}\in S^{(i)}_{3}$to another. We consider this transition to be noiseless. This action is only relevant in Simulation 4, in which we compare the model with or without prior information about the topic. Please note that we only use this factor in Simulation 4.
The tensor $\mathbf{C}^{(i)}=\mathbf{C}_{1}^{(i)}\otimes\mathbf{C}_{2}^{(i)}\otimes\mathbf{C}_{3}^{(i)}$ encodes the priors over the observations that encode preferred outcomes:
  • 1.
    $\mathbf{C}_{1}^{(i)}$, prior preferences on the outcomes $O_{1}^{(i)}$, (e.g. for Level 3 encoding sentences, prior on expected words), which is flat in our simulations;
  • 2.
    $\mathbf{C}_{2}^{(i)}$, prior preferences on the outcomes $O_{2}^{(i)}$ (e.g. for Level 3 encoding sentences, prior on expected word locations), which is flat in our simulations;
  • 3.
    $\mathbf{C}_{3}^{(i)}$, prior preferences on the feedback response $O_{3}^{(i)}$ (e.g. for Level 3 encoding sentences, a higher prior for having classified the sentence according to the ”correct” topic than the ”wrong” topic. Please note that we only use this factor in Simulation 4.
The tensor $\mathbf{D}^{(i)}=\mathbf{D}_{1}^{(i)}\otimes\mathbf{D}_{2}^{(i)}\otimes\mathbf{D}_{3}^{(i)}$ encodes the priors over the hidden states:
  • 1.
    $\mathbf{D}_{1}^{(i)}$, priors over the Content at the corresponding level. It corresponds to the initial distribution of the content variable (e.g. at Level 3, it is the prior over the initial sentence distribution). For simplicity, here we use flat priors. However, it is possible to use this feature to model the fact that syllabes, words and sentences have different frequencies.
  • 2.
    $\mathbf{D}_{2}^{(i)}$, priors on the Locations. For simplicity, here we assign a very high prior to the first element of the content, following the assumption that people start reading (for example) a word from the first syllable. During reading, after each loop the $\mathbf{D}_{1}^{(3)}$ is initialized to the next word not yet read.
  • 3.
    $\mathbf{D}_{3}^{(i)}$, priors on the Topics. We only use this feature in Simulation 4, in which we show that having prior information about the topic of the sentence can speed up its reading. In all the other simulations, we set flat priors over Topics.
    The matrix $\mathbf{E}^{(i)}$ encodes the priors over the policies.
1.4 Saccade selection
Saccade selection arises from a competition among policies spanning all hierarchical levels, determining the location of the next word in the sentence (Level 3), syllable in the word (Level 2), and letter in the syllable (Level 1). The number of policies at each level varies depending on the simulation. For example, when reading 8-word sentences in Simulation 2, the model uses 8 policies to jump among words, 4 policies to jump among syllables and 5 policies to jump among letters.
Policy selection considers the Expected Free Energy (EFE) described in (4). In our simulations, we employ flat prior preferences on Content and Location (denoted as $\mathbf{C}^{(i)}_{1}$ and $\mathbf{C}^{(i)}_{2})$, respectively), resulting in policy selection that solely weighs the information gain associated with the following word, syllable, or letter. However, in Simulation 4, policy selection also depends on the greater preference $\mathbf{C}^{(i)}_{3})$ for receiving a ”correct” than a ”wrong” feedback response (we remind that this factor is absent in the other simulations).
Furthermore, policy selection considers the prior distribution over policies (denoted as $\mathbf{E}^{(i)}$). Our simulations adopt flat priors for Level 1 ($\mathbf{E}^{(1)}$) and Level 2 ($\mathbf{E}^{(2)}$), signifying no inherent constraints on transitioning between letters within a syllable or between syllables within a word. To discourage overly long saccades, a finite moving window for favored transitions among word locations is set for Level 3 ($\mathbf{E}^{(3)}$). This is achieved through a Poisson distribution $Pois(\lambda)$, where $\lambda$ equals 6 (as displayed in Figure 2, depicted by red bars). Additionally, the probability of reading the same word consecutively is set to zero, resulting in the normalization of the distribution. This culminates in the distribution depicted by the black bars in Figure 2. The chosen distribution enables the creation of a finite window to determine probabilities related to the content under consideration. Iterating this window at each step enables the processing of sentences of arbitrary length.
Figure 2: Priors on the policies that make transitions between word locations at Level 3 of the generative model. We first set the prior as a Poisson distribution with $\lambda=6$ (red bars), centered on the current location (i.e., the value of the maximum of the distribution is in 0). We next set the probability of reading the same word two consecutive times to zero and re-normalize the distribution, leading the the distribution shown by the black bars.
Furthermore, Figure 3 presents a schematic illustration of the textual elements inferred by the model while reading the sentence ”Active inference model of eye movements” as depicted in Figure 1. The diagram distinguishes between two factors of the generative model – Content and Location – illustrated by the left and right trees, respectively. The left (content) tree in Figure 3 demonstrates that the model infers the first word ”active” at Level 2 by examining the letter ”a”. This, in turn, leads to the inference of the syllable ”ac” at Level 1, followed by the letters ”t” and ”e”, enabling the inference of the syllable ”tive” at Level 2. The right (location) tree in Figure 3 shows the respective locations of these three letters ”a”, ”t”, and ”e” within their syllables at Level 1, and the positions of ”ac” and ”tive” within the word at Level 2. This process continues for the subsequent words (excluding ”of” and ”eye”, which are skipped during reading, as seen in Figure 3), until the entire sentence is recognized.
Figure 3: Graphical explanation of the model’s reading process for the example sentence ”Active inference model of eye movements” introduced in Figure 1. The figure illustrates the distinction between two factors of the generative model: content (left tree) and location (right tree). Refer to the main text for a detailed explanation.
2 Results
2.1 Simulation 1: Reading single words
Figure 4: Results of Simulations 1 and 2: Probabilities assigned to the correct words. (a) Simulation 1: Probabilities assigned by the Control model (CM) and the Dyslexic model (DM) while reading 4-letter and 8-letter words. (b) Simulation 2: Probabilities assigned by the two models while reading 4-word and 8-word sentences. In this and the subsequent boxplots, the horizontal line represents the mean $\mu$. The edges of the rectangle represent $\mu\pm\sigma$, where $\sigma$ is the standard deviation of the mean. The area within the boxplot above the mean represents the interval between $\mu+\sigma$, while that below the red line represents the interval between $\mu-\sigma$. The vertical black lines limits extend within the range of $\mu\pm 3\sigma$.
In this task, the objective is to read and recognize 100 words consisting of either 4 or 8 letters, selected from a pool of 726 words ranging from 1 to 8 letters, sourced from the BERT dictionary (see Tab. S.14 for the list of words). For ease of evaluation, each word is assigned the same a-priori probability of 1/726 in the model.
The model’s performance is assessed using four metrics: word recognition accuracy, the probability assigned to the correct word, the number of saccades (forward, backward, or total), and their amplitude (i.e., the number of locations between letters). It is important to note that for this specific simulation, only levels 1 and 2 of the model are utilized, while Level 3 is not included.
Previous research has demonstrated that individuals with dyslexia exhibit a reading style characterized by more fragmented and laborious patterns of eye movement, with a higher number of shorter saccades when reading words, and these difficulties are more pronounced with longer words (Zoccolotti et al., 2005; MacKeben et al., 2004; Franzen et al., 2021). To assess whether our model accurately replicates these findings, we compare two versions: (i) a Control model (CM) with no noise in its transition functions, allowing it to adequately consider the prior context during reading, and (ii) a Dyslexic model (DM), where we introduce noise in the transition functions of both the first (syllable) level and the second (word) levels, in line with the hypothesis that dyslexics struggle to integrate prior linguistic context (Jaffe-Dax et al., 2015).
The CM demonstrates perfect accuracy (100%), whereas the DM exhibits slightly lower accuracy for both 4-letter (99%) and 8-letter words (97%). Additionally, the CM assigns a 100% probability to the correct words, whereas the DM assigns a significantly lower probability to the correct word, especially for longer words. Please refer to Figure 4(a) and Table S.1 for detailed results.
Moreover, the DM exhibits a significantly higher number of total saccades (Figure 5(a)) and backward saccades (Figure 5(b)) while reading both 4- and 8-letter words compared to the CM. Detailed statistical comparisons are available in Tab. S.2 and S.3. To delve deeper into the impact of noise on eye movement dynamics during word reading, we created two variants of the DM, one with noise in the transition functions at the syllable level (DM - Noise on Level 1) and another with noise at the word level (DM - Noise on Level 2). This analysis reveals that noise at the first (syllable) level has a more significant effect than noise at the second (word) level.
Figure 5: Simulation 1 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-letter and 8-letter words. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at both hierarchical levels, and two versions of the DM, with noise only at the level of syllables (DM - Noise on Level 1) or words (DM - Noise on Level 2).
In addition, the DM exhibits significantly shorter forward saccades (Figure 6(a)) but longer backward saccades (Figure 6(b)) while reading both 4- and 8-letter words in comparison to the CM. Detailed statistical comparisons are available in Table S.4 and in Table S.5.
Figure 6: Simulation 1 results: Amplitude of forward (a) and backward (b) saccades while reading 4-letter and 8-letter words. The figure compares two models: the Control Model (CM) with no noise, and the Dyslexic Model (DM) with noise at both hierarchical levels.
In summary, our model successfully reproduces a wide range of empirical findings regarding how dyslexic individuals read single words, encompassing accuracy, reaction times, and eye movements. The lower probability assigned to words by the DM compared to the CM, as depicted in Figure 4(a), closely mirrors the results reported by (Zoccolotti et al., 2005, Fig. 1(B)). The accuracy results (Table S.1) are also consistent with the finding that dyslexic individuals fail to correctly read a small percentage of single words. Moreover, our simulations reveal that dyslexic individuals read significantly slower than controls (Zoccolotti et al., 2005; O’Brien et al., 2005, Fig. 1(A)). This slower reading arises from the DM making more saccades than the CM, as illustrated in Figure 5, with reading time being proportional to the number of saccades.
Furthermore, the results displayed in Figure 6(a) closely match the finding that dyslexic individuals make shorter forward saccades compared to controls, particularly evident for longer words (MacKeben et al., 2004, Fig. 4). Lastly, the results presented in Figure 6(b) closely resemble the finding that dyslexic individuals make significantly more and larger backward saccades than controls during reading (Prado et al., 2007, Fig. 5). Collectively, these outcomes demonstrate that our model effectively captures possible aspects of reading skill differences and/or development, providing valuable insights into the underlying mechanisms of reading disorders like dyslexia.
Significantly, our model offers a mechanistic understanding for all these observed findings, including the fragmented reading style of dyslexics. The DM’s shorter forward saccades stem from its poor contextual memory, making it less capable of predicting the next word efficiently. Consequently, it requires more saccades and time to read a text. Additionally, the increased need for backward saccades arises because the DM occasionally has to backtrack in the text to retrieve lost context. Moreover, our model readily explains why dyslexic reading impairments are more pronounced for longer words: noise in the transition function(s) accumulates over time, making reading longer words progressively more challenging. This comprehensive mechanistic explanation underscores the utility of our model in shedding light on the underlying cognitive processes contributing to dyslexia and its impact on reading behavior.
2.2 Simulation 2: Reading Sentences
This task involves reading and recognizing 100 sentences composed of 4 or 8 words, with each word consisting of 1 to 4 syllables, generated by BERT (Devlin et al., 2019). The sentences with 4 words have an average length of $25.96\pm 1.96$ letters within the range of $[21,31]$. Similarly, sentences with 8 words have an average length of $50.24\pm 1.64$ letters, in the interval of $[47,55]$ (see Table S.13). To add complexity, we carefully selected sentences with a substantial word overlap (Table S.15). For this task, we utilize the comprehensive model depicted in Figure 1, encompassing three hierarchical levels.
Previous studies have shown that proficient readers can scan lines of text using only a few saccades, while dyslexic individuals exhibit an increased number of shorter saccades (Hutzler and Wimmer, 2004; Rayner, 1983; De Luca et al., 1999) and fewer word skipping occurrences (Bucci et al., 2008; Jainta and Kapoula, 2011; Hawelka et al., 2010) compared to controls. Additionally, their performance decrease and increased number of saccades are influenced by the number of words in the text (see (De Luca et al., 1999), Fig. 2). To verify whether our model accurately replicates these findings, we compare a Control model (CM) with no noise and a Dyslexic model (DM) with noise introduced in the transition functions of all three hierarchical levels.
Figure 7: Simulation 2 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at both hierarchical levels, and three alternative versions of the Dyslexic Model, where we added noise only at the level of syllables (DM - Noise on Level 1), words (DM - Noise on Level 2), or sentences (DM - Noise on Level 3).
Consistent with the empirical findings, the DM assigns significantly lower probabilities to the correct sentences compared to the CM, and this effect is more pronounced for longer sentences (Figure 4(b) and Table  S.6). Additionally, the DM exhibits a significantly higher number of total saccades (Figure 7(a-c)) and backward saccades (Figure 7(b-d)) than the CM during sentence reading. When comparing different variants of the DM with noise introduced at various levels, we find that noise at the level of sentences has the most significant impact on impairing the reading performance.
Moreover, the DM exhibits significantly shorter forward saccades (Figure 8(a)) and longer backward saccades (Figure 8(b)) when reading both 4-word and 8-word sentences, in comparison to the CM. Detailed statistical comparisons can be found in Table S.9 and in Table S.10.
Collectively, these results build upon those observed for single words (Simulation 1) and successfully replicate empirical findings on dyslexic individuals’ reading of sentences. The trends illustrated in Figure 7(a) closely resemble the outcomes reported (De Luca et al., 1999, Fig. 3(A)), which demonstrates that dyslexics make more saccades compared to controls while reading single lines of text (equivalent to reading 8-word sentences in Figure 7(a)). Moreover, the number of saccades exhibited by the DM in our simulations aligns remarkably well with the numerical outcomes from the same study. Similarly, the outcomes in Figure 8(a) closely resemble the findings depicted in (De Luca et al., 1999, Fig. 3(B)), revealing that dyslexics make shorter forward saccades than controls. In this context, the amplitudes of both dyslexics and the DM model closely correspond.
Figure 8: Results of Simulation 2: amplitude of forward (a) and backward (b) saccades when reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise and the Dyslexic Model (DM) with noise at all the hierarchical levels.
In summary, our results effectively reproduce the observation that dyslexic individuals make a greater number of shorter saccades while reading sentences, and their reading performance is dependent on the number of words in the sentence (refer to (De Luca et al., 1999, Fig.2)). The causes of these impairments are akin to those discussed for single words (Simulation 1), but their effects become more pronounced as sentence comprehension necessitates the integration of information over more extended periods. For instance, the high occurrence of ample backward saccades is a result of the DM frequently losing context and subsequently needing to backtrack to the initial words of a sentence to regain it. This further clarifies why noise at the sentence level substantially impairs this task (Figure 7).
2.3 Simulation 3. Reading novel words and sentences
Up to this point, our simulations involved reading known words and sentences, already present in the BERT vocabulary. Hence, the model could confidently assign probabilities at the second and third layers. Nevertheless, our model can also read unknown words and sentences, despite its inability to assign them a prior probability. When confronted with unknown words or sentences, the model reads syllable-by-syllable (or word-by-word) and simultaneously acknowledges uncertainty at the second (or third) level. This mode aligns with the nonlexical (or sublexical) route in dual-route theories of reading (Coltheart, 2005).
To demonstrate this capability, we simulate reading the last four words of the sentence This paper is also framed in an offbeat manner and compare two versions of the model: one in which all the words exist in the model’s vocabulary (Figure 9(a)) and another where the word offbeat is absent from the vocabulary (Figure 9(b)). In the version with all known words, the model behaves similarly to the previous simulations, accurately recognizing all the words and the sentence. This is evident in Figure 9(a), where probabilities at the second and third levels quickly converge to one after a few iterations.
Conversely, the version without knowledge of the word offbeat repeatedly makes saccades to the syllables of this word (at Level 1) until it reaches the maximum number of iterations set for this simulation. It correctly recognizes each syllable and can read the word aloud, potentially adding it to its vocabulary (Coltheart, 2005). However, it does not recognize the word (at Level 2) or the sentence (at Level 3). This is evident in Figure 9(b), where probabilities at the second and third levels never exceed 0.5.
The example in Fig. 9 illustrates the distinction between reading a known versus a novel word and the termination conditions for inference at each hierarchical level. The inference can terminate in two ways: first, when the model reaches a threshold of confidence (entropy of hidden states, $\chi^{(i)}$) about the to-be-recognized syllable, word, or sentence. This happens when the model successfully infers a known syllable, word, or sentence.
Second, the inference can terminate when it reaches a maximum number of iterations $K_{\max}^{(i)}$ at each $i$-th level. This sets the maximum number of times Level 1 can jump between letters to recognize a syllable, the maximum number of times Level 2 can jump between syllables to recognize a word, or the maximum number of times Level 3 can jump between words to recognize a sentence. In our simulations, for illustrative purposes, we set $K_{\max}^{(i)}$ to very high values ($K_{\max}^{(3)}=7$ and $K_{\max}^{(2)}=K_{\max}^{(1)}=6$) but these values can be fine tuned to fit individual participants or participant groups.
As shown in Figure 9, Level 1 successfully infers all the syllables within 3 iterations and never reaches $K_{\max}^{(1)}$. However, Level 2 fails to recognize the unknown word offbeat: it continues jumping between the syllables off and beat until it reaches the maximum number of allowed iterations (6 iterations, indicated in the grey panel). When the Level 2 inference of the current word is halted, a low confidence message is reported to the level above and the inference continues with the next words (that are successfully recognized). Similarly, Level 3 fails to infer the unknown sentence after the maximum number of iterations (7 iterations), despite correctly recognizing most of the words. At the end of the simulation, the model identifies that the cause of failure is at Level 2, as Level 1 correctly recognized all the syllables. Consequently, the model reads the unknown word syllable-by-syllable, i.e. using the nonlexical route in dual route theories of reading (Coltheart, 2005).
Figure 9: Simulation 3 results: Reading known vs. unknown words. Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time, while reading (a) known words or (b) the unknown word offbeat. The blue circles at the bottom level indicate saccades. The figure format is consistent with Figure 1(d).
2.4 Simulation 4: Knowing the topic improves sentence recognition
In this section, we showcase another capability of the model that has not been discussed yet. The model can be enriched with priors regarding the linguistic topic at each hierarchical level, such as the topic of the sentence at the third level. This prior knowledge makes certain sentences a-priori more (or less) likely, thus expediting the recognition process if the context aligns correctly.
To illustrate this, we conduct a task involving the recognition of 100 sentences, each containing 9 words, generated by BERT. These sentences comprise an average of $69.55\pm 5.02$ letters, ranging between 61 and 82 characters (see Table S.13). We allocate one-third of the sentences to each of the three topics of the European Research Council (ERC): Physical Sciences and Engineering (PE), Life Sciences (LS), and Social Sciences and Humanities (SH).
We compare the performance of two versions of the model: one with a flat prior (33.3%) regarding the sentence being read, and another with an informative prior that assigns a probability of 100% to the correct topic. As depicted in Figure  10, the model with an informative prior requires significantly fewer total saccades and backward saccades than the model with a flat prior (see Table S.11 and Table S.12 for statistical comparisons). This advantage arises because knowing the topic allows the model to predict certain words with high confidence, enabling it to skip unnecessary saccades during the reading process. Intriguingly, despite the model with an informative prior skipping several words, it does not necessitate more backward (corrective) saccades compared to the model with a flat prior. This indicates that the prior information promotes a better speed-accuracy trade-off.
Figure 10: Simulation 4 results: How prior information affects the reading process. The figure displays the total saccades (a) and backward saccades (b) of two model versions: one with a flat prior about the topic of the text (flat prior) and another with certain knowledge of the topic (informative prior). The comparison highlights the impact of prior information on the number and nature of saccades during reading.
3 Discussion
In this paper, we presented a novel hierarchical active inference model of reading, which combines the robust generative abilities of large language models (Devlin et al., 2019; Radford et al., 2019) with the inferential and information-gathering capacities of active inference (Friston, 2010; Parr et al., 2022).
Large language models, like BERT (Devlin et al., 2019) and GPT (Generative Pre-trained Transformer) (Radford et al., 2019), have demonstrated remarkable performance in various natural language processing tasks, including question answering, language translation, text classification, and sentence prediction. These models can learn essential linguistic structures without external supervision, making them valuable tools to explore language processing in the brain (Manning et al., 2020; Goldstein et al., 2022). However, they lack the active strategies employed by human readers, such as the ability to make saccades selectively on the most informative parts of the text (Ferro et al., 2010; Norris, 2006).
To capture active reading, we integrate BERT (Devlin et al., 2019) into a hierarchical active inference framework (Friston, 2010; Parr et al., 2022). Active inference is a theoretical framework that describes how the brain employs probabilistic inference over a generative model to effectively sample sensory inputs, minimizing variational free energy or prediction errors. For modeling reading, we utilize a hierarchical generative model, aligning with evidence suggesting that language processing in the brain operates through hierarchically organized predictions and prediction errors (Schmitt et al., 2021; Caucheteux et al., 2023; Heilbron et al., 2022).
By placing BERT at the top of the hierarchy, we ensure accurate next-word predictions and the model’s capability to tackle real-world reading tasks. Beyond next-word prediction, our hierarchical model facilitates both inferring the content being read and predicting future elements across different linguistic levels, such as syllables, words, and sentences. Our simulation results demonstrate the model’s perfect accuracy in inferring words (Simulation 1) and sentences (Simulation 2) during reading. Additionally, this hierarchical structure enables the model to read novel words not present in BERT’s vocabulary (Simulation 3), assembling them syllable-by-syllable, representing the nonlexical (or sublexical) route in dual route theories of reading (Coltheart, 2005). Moreover, the model can incorporate prior knowledge, like the topic of the sentence, to expedite the reading process (Simulation 4). Significantly, the model offers interpretability through its alignment of distinct levels with established linguistic constituents (sentences, words, syllables), which could correspond to distributed neural populations or cell assemblies within the brain (Heilbron et al., 2022; Pulvermüller, 1999). While delving into the neural mechanisms that underlie this proposed framework extends beyond the confines of this article, future studies could encompass the simulation of neuronal dynamics during the inference process, in terms of variational free energy minimization (Friston et al., 2017; Isomura et al., 2022).
Using active inference, our model characterizes reading as an active, hypothesis testing process, allowing us to simulate eye movements during reading. The core idea is that saccades are directed to the most informative parts of the text, enabling the model to test its predictions and reduce uncertainty about the text being read (Ferro et al., 2010; Norris, 2006; Donnarumma et al., 2017; Friston et al., 2012). Remarkably, this epistemic objective, aimed at uncertainty reduction, emerges naturally from active inference, as explained in Section 1. In summary, our model offers a comprehensive account of prediction-based written text processing, involving hierarchical inference of the content being read, prediction of upcoming textual elements, and active testing of predictions through saccades, which, in turn, inform further inference.
Importantly, our model also enables the modeling of abnormal eye movements observed in reading disorders, such as dyslexia (Maisog et al., 2008; Horowitz-Kraus et al., 2014; Price, 2012). Our simulations replicate the empirical findings that dyslexics, compared to controls, exhibit fragmented text processing, characterized by an increased number of shorter saccades, both while reading single words (Simulation 1) and sentences (Simulation 2) (Zoccolotti et al., 2005; De Luca et al., 1999; Hutzler and Wimmer, 2004; Rayner, 1983). Interestingly, our model attributes these reading deficits to a straightforward mechanism: an attenuation of the role of prior information during inference (Jaffe-Dax et al., 2015). By introducing noise in the model’s transition functions, we impair its working memory of previous inference phases, forcing it to repeatedly gather evidence. Considering the cumulative effect of noise on the inference process, this mechanism accounts for why perceptual deficits in dyslexics become more pronounced for longer words and sentences (Zoccolotti et al., 2005; De Luca et al., 1999). To summarize, the model provides a mechanistic explanation of reading disorders in dyslexia, aligning well with prediction-based theories of language processing and reading.
However, it is crucial to note that dyslexia is highly heterogeneous, and its underlying causes are still heavily debated (Perry et al., 2019; Zorzi et al., 2012; Ramus, 2003; Stein and Walsh, 1997; Tallal, 2004). While the attenuation of priors (Jaffe-Dax et al., 2015) explored in this article may not account for all aspects of dyslexia, our model is easily extendable to incorporate other (non-alternative) mechanisms that could contribute to reading impairments. For instance, various proposals point to disorders of low-level information sampling and attention shifting (Hari et al., 2001), or a visual attention span disorder characterized by a reduction in the number of letters that can be processed in parallel (Prado et al., 2007). These proposals (and others) can be readily integrated into the model by modifying the likelihood function that maps letters to syllables, as described in Section 1. Future research could utilize different variants of the presented model (e.g., with different levels of noise in the likelihood function and/or the transition function) to compare alternative proposals side-by-side, shedding light on the mechanisms of dyslexia and other reading difficulties, and informing the development of effective interventions to improve reading skills.
Acknowledgments
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 945539 (Human Brain Project SGA3) and No. 952215 (TAILOR) to GP; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to GP; the PNRR MUR projects PE0000013-FAIR and IR0000011–EBRAINS-Italy to GP; the MUR PRIN2020 project “Free energy principle and the brain: neuronal and phylogenetic mechanisms of Bayesian inference” - Grant N. 2020529PCP to FD. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Supplementary materials
Integrating Large Language Models
and Active Inference
to understand eye movements
in Reading and Dyslexia
Francesco Donnarumma
Mirco Frosolone
Giovanni Pezzulo
Institute of Cognitive Sciences and Technologies
National Research Council
Via San Martino della Battaglia, 44
00185 Rome Italy
Corresponding author: Giovanni Pezzulo E-mail: giovanni.pezzulo@istc.cnr.it
Supplementary info
In this document, we provide the tables of statistics from the simulations presented in the main article (found in Supplementary Tables section) and the Datasets of sentences and words used in the simulation (found in Datasets section).
Supplementary tables
Below are the supplementary tables presenting statistics related to the simulations discussed in the main manuscript.
Supplementary tables for Simulation 1
Table S.1 displays the accuracy and probability assigned to the correct words. Table S.2 shows the T-tests on total saccades and Table S.3 the backward saccades. Table S.4 and Table S.5 presents the T-tests for the amplitude of forward and backward saccades respectively. The $p$-values lower to $0.05$ are indicated in bold.
Table S.1: Simulation 1: Reading words of 4 or 8 letters. The table presents the accuracy and the probability assigned to the correct words by 4 different models: the Control model (CM), the Dyslexic model (DM) variant with noise at Level 1 only, the Dyslexic model (DM) variant with noise at Level 2 only, and the Dyslexic model (DM) with noise at both levels. See Figure 4 and the main text for details.
4 - letter 8 - letter
Accuracy Probability Accuracy Probability
Control Model $100$ $1.00\pm 1.03\cdot 10^{-14}$ $100$ $1.00\pm 7.85\cdot 10^{-15}$
DM -Noise on Level 1 $100$ $1.00\pm 3.32\cdot 10^{-04}$ $100$ $0.99\pm 3.82\cdot 10^{-04}$
DM -Noise on Level 2 $99$ $0.99\pm 3.73\cdot 10^{-15}$ 97 $0.97\pm 1.38\cdot 10^{-02}$
Dyslexic Model $99$ $0.99\pm 3.32\cdot 10^{-04}$ $97$ $0.97\pm 1.15\cdot 10^{-02}$
Table S.2: Simulation 1. T-tests for simulation of 4-letter and 8-letter word reading. The table displays the results of T-tests conducted on the total number of saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, and the Dyslexic model (DM) with noise at both levels. See Figure 5 and the main text for details.
Total number of saccades
4 - letter 8 - letter
DM - Noise Level 1 $F[1,198]=17.6,\mathbf{{\bf\it p}\approx 4.1\cdot 10^{-05}}$ $F[1,198]=12.7,\mathbf{{\bf\it p}\approx 4.5\cdot 10^{-04}}$
DM - Noise Level 2 $F[1,196]=4.4,\mathbf{{\bf\it p}\approx 3.6\cdot 10^{-02}}$ $F[1,190]=5.2,\mathbf{{\bf\it p}\approx 2.4\cdot 10^{-02}}$
Dyslexic Model $F[1,196]=16.7,\mathbf{{\bf\it p}\approx 6.4\cdot 10^{-05}}$ $F[1,192]=10.7,\mathbf{{\bf\it p}\approx 1.3\cdot 10^{-03}}$
Table S.3: Simulation 1. T-tests for simulation of 4-letter and 8-letter word reading. The table displays the results of T-tests conducted on the number of backward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, and the Dyslexic model (DM) with noise at both levels. See Figure 5 and the main text for details.
Number of backward saccades
4 - letter 8 - letter
DM - Noise Level 1 $F[1,198]=21.5,\mathbf{{\bf\it p}\approx 6.5\cdot 10^{-06}}$ $F[1,198]=17.5,\mathbf{{\bf\it p}\approx 4.4\cdot 10^{-05}}$
DM - Noise Level 2 $F[1,198]=4.8,\mathbf{{\bf\it p}\approx 3.0\cdot 10^{-02}}$ $F[1,190]=5.7,\mathbf{{\bf\it p}\approx 1.8\cdot 10^{-02}}$
Dyslexic Model $F[1,196]=21.7,\mathbf{{\bf\it p}\approx 5.8\cdot 10^{-06}}$ $F[1,192]=13.3,\mathbf{{\bf\it p}\approx 3.4\cdot 10^{-04}}$
Table S.4: Simulation 1. T-tests for simulation of 4-letter and 8-letter word reading. The table displays the results of T-tests conducted on the amplitude of forward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, and the Dyslexic model (DM) with noise at both levels. See Figure 6 and the main text for details.
Amplitude of forward saccades
4 - letter 8 - letter
Dyslexic Model $F[1,844]=9.4,\mathbf{{\bf\it p}\approx 2.2\cdot 10^{-03}}$ $F[1,678]=46.5,\mathbf{{\bf\it p}\approx 2.0\cdot 10^{-11}}$
Table S.5: Simulation 1. T-tests for simulation of 4-letter and 8-letter word reading. The table displays the results of T-tests conducted on the amplitude of + backward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, and the Dyslexic model (DM) with noise at both levels. See Figure 6 and the main text for details.
Amplitude of backward saccades
4 - letter 8 - letter
Dyslexic Model $F[1,321]=133.5,\mathbf{{\bf\it p}\approx 4.6\cdot 10^{-26}}$ $F[1,312]=24.5,\mathbf{{\bf\it p}\approx 1.2\cdot 10^{-06}}$
Supplementary tables for Simulation 2
Table S.6 display the accuracy and probability assigned to the correct words. Table S.7 and Table S.8 shows the T-tests on total saccades and back saccades. Table S.9 presents the T-tests for amplitude of forward and Table S.10 the T-test for amplitude of backward saccades respectively. The $p$-values lower to $0.05$ are indicated in bold.
Table S.6: Simulation 2: Reading sentences of 4 or 8 words. The table presents the accuracy and the probability assigned to the correct words by 4 different models: the Control model (CM), the Dyslexic model (DM) variant with noise at Level 1 only, the Dyslexic model (DM) variant with noise at Level 2 only, the Dyslexic model (DM) variant with noise at Level 3 only, and the Dyslexic model (DM) with noise at all levels. See Figure 4 and the main text for details.
4 - word 8 - word
Accuracy Probability Accuracy Probability
Control Model $100$ $1.00\pm 3.66\cdot 10^{-16}$ $100$ $1.00\pm 3.21\cdot 10^{-12}$
DM - Noise on Level 1 $100$ $1.00\pm 3.64\cdot 10^{-06}$ $100$ $1.00\pm 5.69\cdot 10^{-08}$
DM - Noise on Level 2 $100$ $1.00\pm 3.40\cdot 10^{-07}$ $100$ $1.00\pm 3.21\cdot 10^{-12}$
DM - Noise on Level 3 $97$ $0.97\pm 4.25\cdot 10^{-03}$ $93$ $0.93\pm 1.04\cdot 10^{-02}$
Dyslexic Model $96$ $0.96\pm 4.41\cdot 10^{-03}$ $93$ $0.93\pm 1.04\cdot 10^{-02}$
Table S.7: Simulation 2. T-tests for simulation of 4-word and 8-word sentence reading. The table displays the results of T-tests conducted on the total number of saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 7 and the main text for details.
Total number of saccades
4 - word 8 - word
DM - Noise Level 1 $F[1,198]=0.02,~{}p\approx 0.90$ $F[1,198]=0.01,~{}p\approx 0.92$
DM - Noise Level 2 $F[1,198]=0.97,~{}p\approx 0.453$ $F[1,198]=0.451,~{}~{}p\approx 0.58$
DM - Noise Level 3 $F[1,198]=66.8,\mathbf{{\bf\it p}\approx 3.5\cdot 10^{-14}}$ $F[1,196]=105.9,\mathbf{{\bf\it p}\approx 3.9\cdot 10^{-20}}$
Dyslexic Model $F[1,198]=75.5,\mathbf{{\bf\it p}\approx 1.4\cdot 10^{-15}}$ $F[1,196]=113.6,\mathbf{{\bf\it p}\approx 3.2\cdot 10^{-21}}$
Table S.8: Simulation 2. T-tests for simulation of 4-word and 8-word sentence reading. The table displays the results of T-tests conducted on the number of backward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 7 and the main text for details.
Number of backward saccades
4 - word 8 - word
DM - Noise Level 1 $F[1,198]=0.451,~{}p\approx 0.52$ $F[1,198]=0.82,~{}p\approx 0.457$
DM - Noise Level 2 $F[1,198]=0.01,~{}p\approx 0.90$ $F[1,198]=1.86,~{}p\approx 0.17$
DM - Noise Level 3 $F[1,198]=44.7,\mathbf{{\bf\it p}\approx 2.3\cdot 10^{-10}}$ $F[1,196]=119.4,\mathbf{{\bf\it p}\approx 5.1\cdot 10^{-22}}$
Dyslexic Model $F[1,198]=53.2,\mathbf{{\bf\it p}\approx 7.1\cdot 10^{-12}}$ $F[1,196]=117.7,\mathbf{{\bf\it p}\approx 8.8\cdot 10^{-22}}$
Table S.9: Simulation 2. T-tests for simulation of 4-word and 8-word sentence reading. The table displays the results of T-tests conducted on the amplitude of forward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 8 and the main text for details.
Amplitude of forward saccades
4 - word 8 - word
Dyslexic Model $F[1,878]=18.0,\mathbf{{\bf\it p}\approx 2.5\cdot 10^{-05}}$ $F[1,1238]=10.9,\mathbf{{\bf\it p}\approx 9.8\cdot 10^{-04}}$
Table S.10: Simulation 2. T-tests for simulation of 4-word and 8-word sentence reading. The table displays the results of T-tests conducted on the amplitude of backward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 8 and the main text for details.
Amplitude of backward saccades
4 - word 8 - word
Dyslexic Model $F[1,318]=14.7,\mathbf{{\bf\it p}\approx 1.5\cdot 10^{-04}}$ $F[1,350]=42.5,\mathbf{{\bf\it p}\approx 2.5\cdot 10^{-10}}$
Simulation 4
Table S.11 and Table S.12 presents T-tests on total number of saccades and the number backward saccades when the model is augmented with the topic of the sentence at Level 3 of the hierarchy. The $p$-values lower to $0.05$ are indicated in bold.
Table S.11: Simulation 4. T-tests for simulation of a 9-word sentence reading, belonging to different topics. The table displays the results of T-tests conducted on the total number of saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 10 and the main text for details.
Total number of saccades
Flat Priors VS Informative Priors $F[1,198]=4.07,\mathbf{{\bf\it p}\approx 4.1\cdot 10^{-03}}$
Table S.12: Simulation 4. T-tests for simulation of a 9-word sentence reading, belonging to different topics. The table displays the results of T-tests conducted on the number of backward saccades for different models with respect to the Control model (CM): the Dyslexic model (DM) variant with noise at Level 1 (syllable) only, the Dyslexic model (DM) variant with noise at Level 2 (word) only, the Dyslexic model (DM) variant with noise at Level 3 (sentence) only, and the Dyslexic model (DM) with noise at all levels. See Figure 10 and the main text for details.
Number of backward saccades
Flat Priors VS Informative Priors $F[1,198]=2.57,~{}~{}p\approx 0.11$
Datasets
In this section, we present the datasets utilized in our simulations. Information regarding content length in the datasets can be found in Table S.13. Furthermore, Table S.14 displays the dictionary of words used in Simulation 1, while Table S.15 provides insight into the sentences used in Simulation 2. Additionally, Simulation 4’s sentences and corresponding labels are showcased in Table S.16.
Table S.13: This table reports the mean number of characters, the standard deviation, minimum and maximum number of characters that compose the relative set of sentences used in the simulations.
Mean
Standard
Deviation
Min Max
Simulation 2 (4 - word) $25.96$ $1.96$ $21$ $31$
Simulation 2 (8 - word) $50.24$ $1.64$ $47$ $55$
Simulation 4 $69.55$ $5.02$ $61$ $82$
Table S.14: Dictionary used in Simulation 1 (words extracted from BERT Dictionary)
1 - letter 2- letter 3- letter 4- letter 5- letter 6- letter 7- letter 8- letter
A AC AMC ARES AGREE ACTING AIRPORT ACQUIRED
B AL ARM ARTS ANZAC ADVENT ALCOHOL ACTUALLY
C AM ATA ASKS ARDEN AIRMEN ANGRILY ADVANCED
D AR AUS BEDS ARIEL ANGLIA ARCADIA ALPHABET
E AW BAO BORE AZURE ARMAND ASSURED AMERICAN
F AX BEG BRET BARED AUSTEN AVENUES APPLETON
G BF BEN BUSY BEGUN BECKER BALANCE ARGUMENT
H BI BIT BUTT BIRDS BEIRUT BENGALS ARRANGED
I BK BOP CALE BLAND BERMAN BERWICK ATLANTIC
J BO BUY CARL BUICK BIGGER BOUNCED BALANCED
K BR CAB COAL BULGE BISHOP BURUNDI BASILICA
L BT CAD COLA BUTCH BUMPER CAREERS BENEDICT
M CA CAI COOL CELLO BURDEN CEILING BOTSWANA
N CB CEO CUBS CLEAR CANTOR CHANCEL BOULDERS
O CH COE DIRT CLUBS CAUSAL CLICKED BREACHED
P CL CPU DOIN CODES CEREAL CLIENTS BREAKING
Q CM DEX DOVE COSTS CESARE COMMITS CADILLAC
R CN DIG DRIP CRIED CESSNA CROOKED CAMPBELL
S CO DIP DRUM CURLY CHICKS DERRICK CARRYING
T CP EKI ELLE DANNY CLOCKS DESCENT CATALINA
U CU EPA EMIR DATED COINED DESPAIR COLISEUM
V ED ETA EMMA DEBUT DECKER DEVISED COLUMBIA
W EH FIG FANS DISCS DISMAY EDUARDO CONNECTS
X EL FLU FEES DOUBT DRINKS EMBASSY CONVINCE
Y EM FUN FIJI DUCKS EXEMPT EMERSON CUMMINGS
Z EU HAN FINS DUMMY FACADE EMPLOYS CURRENCY
FI HBO FLOP ECOLE FINLEY ENEMIES CYCLONES
FL HEM FLUX ELIZA GENTRY ESCAPES CYLINDER
FT HIT FRAN ESSEN HAILEY EXHAUST DARKENED
FU HOP GIFT FEMME HANNAH FAINTLY DECEASED
FX HUE GIGS FISTS HASSAN FLEDGED DECIDING
GE IBN GLOW FLOOR HATRED FORSTER DISABLED
GI ICE GRAM FLOYD HELMUT FREEWAY DISASTER
GO ICH GREY GOUGH HOOVER FRONTED DONNELLY
GS ICT HAAS GROWS HUTTON GALILEO EMERITUS
GT ILE HANK GRUNT HYBRID GEOLOGY EMPLOYER
GU ION HILL GUARD INCOME GLIDING EPILOGUE
HA IRA HISS HELLO INPUTS HAMBURG EREBIDAE
HC KIM HUEY HIRES INTERN HANDLED EVACUATE
HM LAL INDY HORDE JOSEPH HIGHEST EVOLVING
HO LAY JACE JOINT JUAREZ HONNEUR EXAMINER
HU LEG JAIN JOYCE KHYBER HORNETS FABULOUS
IK LES JEAN LATER LANDON HOWEVER FIGHTERS
IN LOU JING LEANS LENGTH HOWLING FRANKLIN
IO LTD KRIS LEAVE LENNON HUNCHED FRICTION
IR MED LAMA LOOSE LIKELY INHALED GESTURES
IX MPH LANG LOVER LILITH IRANIAN GROUPING
JA MPS LETO LUCHA LITTER JEALOUS HOMICIDE
JD MSC LIAR LYDIA MADRAS JUSTINE IMPERIAL
KE MUD MACK MAHAL MANUEL KNIGHTS INCURRED
KG NBA MARS MERGE MARGIN LABELED INITIATE
KN NHL MASK MINSK MARKED LASTING INJURIES
KO NHS MEAT MISSY MCLEAN LEBANON INSANITY
KS NIK MIKA MIXED MOLINA MANNING JUDICIAL
LC NOS MOJO MOMMA MUSEUM MCBRIDE LABRADOR
LI NOV MOOD NAILS MYSELF MORALLY MAGAZINE
LP NPR MUCH NORMA NEWELL MUSCLED MARATHON
MG NYC NEAT OLDER ONIONS NATALIE MEANINGS
MI OFF NINO OUTTA OPENLY NEURONS MOHAMMAD
ML OUT NOSE PAPUA OPPOSE NEVILLE MONSIEUR
MR PAT NRHP PATIO PANZER NIKOLAI MOROCCAN
NA PAU OBOE PIANO PAVING NILSSON MOUNTING
NH PBA OWLS PILES PAYTON NOMINAL MUSHROOM
NI PBS PASS PLUTO PETALS OFFENSE OLYMPIAD
NO PCS PERU PROSE PILOTS ONSTAGE PACKAGES
NS PEI PORN QAEDA PLANES PATCHES PATERSON
NT PEW RAAF RALLY POLITE PEPTIDE PHILLIPS
NZ PHI RAMP REEFS POTION PERFECT POPULACE
OG PHP RAMS REVUE PRAGUE PEUGEOT POUNDING
OH POD RUDD ROACH PRICED PIONEER PREACHER
OL PUN SANK ROUEN PULSES PLANNED PRESTIGE
OP ROE SAXE RUSSO PUPILS PLAYFUL QUARTERS
OS ROM SITU SCOTT RASHID PORTICO RAILROAD
OZ RUM SIZE SHORE REPORT PROTEST RECEIVES
PH RYE SLID SINGH RIPLEY RAPIDLY REJOINED
PO SAO SODA SITED ROBERT READERS SCENARIO
PR SEA SOON SLAMS RUNNER RECITAL SEMINOLE
RC SHI STAY SPRAY SAILOR REDWOOD SHOOTOUT
RR SHU STYX SQUAT SAXONY REISSUE SHOWERED
RY SPP SUFI STADE SCARES ROOSTER SINGULAR
SE SSR SWAM STAFF SCREAM RUINING SOMERSET
SF SUE TEAM STALL SEIZED SECURED SOUTHEND
SM TAI THRU STATE SESAME SILENCE SPECIALS
SQ TEA TODD STERN SEVERN SMASHED SPEEDING
SR TEX TOLD STORM SIGNAL SPELLED STRANDED
TC TIS TOPS STUMP SPEAKS STATION STRIPPED
TU TNA TRIP TOUGH SPINES STEWART SYMBOLIC
TV UFO TUNA TRANS STALLS STRANGE TARGETED
TX VAR UGLY TREAT STARTS STRIKES TAXATION
UP VII USER TREES SURFER SWELLED THROTTLE
UR WAT VEGA VAULT TORINO TIGHTLY TIMELINE
VA WAY VIDA VIJAY TREVOR TOPICAL TOURISTS
VE WOW VISA VISAS TURTLE TURTLES UNCOMMON
VP XII WEEP VOMIT VENDOR ULYSSES VALENTIN
VU XVI WHOA WEARS VERBAL UNITING VIGOROUS
WC YES WIFE WHORE WANTED UNNAMED VILLAINS
XI YET WINS WRAPS WASHED VACANCY WARRIORS
YA YOO WITS WRIST WIRING VISIBLE WEREWOLF
YE ZEE YUKI YEMEN WORTHY WALKING WORRYING
YO ZEV ZOOM YOUTH XAVIER WARTIME ZIMBABWE
Table S.15: Dictionary used in Simulation 2
FAITH CONQUERS ALL OBSTACLES
FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH CONQUERS EACH HURDLE
FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH CONQUERS EACH SETBACK
FAITH CONQUERS EACH HURDLE BUT TRASCENDS ALL MISFORTUNE
FAITH CONQUERS EVERY CHALLENGES
FAITH CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
FAITH CONQUERS EVERY OBSTACLES
FAITH CONQUERS EVERY OBSTACLES AND PREVAILS OVER ADVERSITY
FAITH DEFEATS ALL OBSTACLES
FAITH CONQUERS EVERY OBSTACLES AND SURMOUNTS ALL OBSTACLES
FAITH OVERCOMES ALL HINDRANCES
FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH OVERCOMES ALL OBSTACLES
FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH OVERPOWERS ALL LIMITATIONS
FAITH OVERCOMES ALL HINDRANCES AND TRASCENDS OVER MISFORTUNE
FAITH PREVAILS AGAINST BARRIERS
FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH PREVAILS AGAINST OBSTACLES
FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH PREVAILS OVER EVERY ADVERSITY
FAITH OVERPOWERS ALL LIMITATIONS BUT TRASCENDS ALL MISFORTUNE
FAITH PREVAILS OVER OBSTACLES
FAITH PREVAILS AGAINST BARRIERS AND TRASCENDS OVER MISFORTUNE
FAITH SURMOUNTS ALL OBSTACLES
FAITH PREVAILS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH SURPASSES ALL OBSTACLES
FAITH PREVAILS EVERY ADVERSITY BUT SURMOUNTS ALL OBSTACLES
FAITH SURPASSES ALL OPPOSITION
FAITH PREVAILS OVER ADVERSITY BUT SURMOUNTS ALL OBSTACLES
FAITH TRIUMPHS AGAINST OBSTACLES
FAITH PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH TRIUMPHS OVER CHALLENGES
FAITH SURMOUNTS ALL OBSTACLES AND SURMOUNTS ALL MISFORTUNE
FAITH TRIUMPHS OVER OBSTACLES
FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH VANQUISHES ALL OBSTACLES
FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH VANQUISHES EVERY DIFFICULTY
FAITH SURPASSES ALL OBSTACLES BUT SURMOUNTS ALL ADVERSITY
HOPE CONQUERS ALL LIMITATIONS
FAITH SURPASSES ALL OPPOSITION BUT SURMOUNTS ALL MISFORTUNE
HOPE CONQUERS ALL OBSTACLES
FAITH TRIUMPHS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH CHALLENGES
FAITH TRIUMPHS OVER CHALLENGES AND TRASCENDS OVER MISFORTUNE
HOPE CONQUERS EACH HINDRANCES
FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EACH HURDLE
FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH INHIBITION
FAITH VANQUISHES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH LIMITATIONS
FAITH VANQUISHES EVERY DIFFICULTY BUT TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH SETBACK
HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY ADVERSITY
HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY BARRIERS
HOPE CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
HOPE CONQUERS EVERY CHALLENGES
HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY DIFFICULTY
HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY HURDLE
HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY OBSTACLES
HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY RESTRAINT
HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE DEFEATS ALL OPPOSITION
HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE OVERCOMES ALL CHALLENGES
HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE OVERCOMES EACH BOUNDARY
HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE OVERCOMES EACH HINDRANCES
HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE PREVAILS AGAINST ADVERSITY
HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE PREVAILS AGAINST BARRIERS
LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL ADVERSITY
HOPE PREVAILS AGAINST IMPEDIMENT
LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL MISFORTUNE
HOPE PREVAILS OVER OBSTACLES
LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS ALL MISFORTUNE
HOPE SURMOUNTS ALL DIFFICULTIES
LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS EVERY MISFORTUNE
HOPE SURPASSES ALL OBSTACLES
LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL ADVERSITY
HOPE SURPASSES EVERY CONFINEMENT
LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL MISFORTUNE
HOPE SURPASSES EVERY OBSTACLES
LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS ALL MISFORTUNE
HOPE TRANSCENDS EACH RESTRICTION
LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS EVERY MISFORTUNE
HOPE TRIUMPHS AGAINST OBSTACLES
LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL ADVERSITY
HOPE TRIUMPHS OVER CHALLENGES
LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL MISFORTUNE
HOPE VANQUISHES ALL HINDRANCES
LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS ALL MISFORTUNE
LOVE OVERCOMES EACH HINDRANCES
LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS EVERY MISFORTUNE
LOVE CONQUERS ALL ADVERSITY
LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL BOUNDARIES
LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE CONQUERS ALL DIFFERENCES
LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL OBSTACLES
LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS EVERY ADVERSITY
LOVE CONQUERS ANY OBSTACLES
LOVE CONQUERS EVERY BARRIERS AND PREVAILS OVER ADVERSITY
LOVE CONQUERS EACH OBSTACLES
LOVE CONQUERS EVERY OBSTACLES BUT TRASCENDS ALL MISFORTUNE
LOVE CONQUERS EVERY BARRIERS
LOVE DEFEATS ALL OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE CONQUERS EVERY CHALLENGES
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS ALL ADVERSITY
LOVE CONQUERS EVERY CONFINEMENT
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS ALL MISFORTUNE
LOVE CONQUERS EVERY DIFFICULTY
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS EVERY OBSTACLES
LOVE CONQUERS EVERY HINDRANCES
LOVE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
LOVE CONQUERS EVERY HURDLE
LOVE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE CONQUERS EVERY LIMITATIONS
LOVE OVERCOMES ALL OBSTACLES BUT CONQUERS ALL ADVERSITY
LOVE CONQUERS EVERY OBSTACLES
LOVE OVERCOMES ALL OBSTACLES BUT CONQUERS EVERY ADVERSITY
LOVE CONQUERS EVERY SETBACK
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE DEFEATS ALL OBSTACLES
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS EACH ADVERSITY
LOVE DEFEATS ALL OPPOSITION
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS EVERY ADVERSITY
LOVE OVERCOMES ALL OBSTACLES
LOVE PREVAILS AGAINST CHALLENGES AND PREVAILS OVER ADVERSITY
LOVE PREVAILS AGAINST CHALLENGES
LOVE PREVAILS AGAINST OBSTACLES AND OVERCOMES OVER MISFORTUNE
LOVE PREVAILS AGAINST OBSTACLES
LOVE PREVAILS AGAINST OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE PREVAILS OVER OBSTACLES
LOVE PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE SURMOUNTS ALL OBSTACLES
LOVE SURMOUNTS ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE SURPASSES ALL OBSTACLES
LOVE SURMOUNTS ALL OBSTACLES BUT TRASCENDS ALL MISFORTUNE
LOVE TRIUMPHS AGAINST OBSTACLES
LOVE SURPASSES ALL OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE TRIUMPHS OVER OBSTACLES
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS ALL ADVERSITY
LOVE VANQUISHES ALL OBSTACLES
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS ALL OBSTACLES
UNITY BREAKS THROUGH ALL BARRIERS
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS EACH ADVERSITY
UNITY CONQUERS AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS OVER ADVERSITY
UNITY CONQUERS ALL OBSTACLES
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS OVER OBSTACLES
UNITY CONQUERS DESPITE ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS ALL ADVERSITY
UNITY DEFEATS AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS ALL OBSTACLES
UNITY DEFEATS ALL OPPOSITION
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS EACH ADVERSITY
UNITY OVERCOMES AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS OVER ADVERSITY
UNITY OVERCOMES ALL CHALLENGES
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS OVER OBSTACLES
UNITY OVERPOWERS ALL RESISTANCE
LOVE TRIUMPHS OVER OBSTACLES AND PREVAILS OVER ADVERSITY
UNITY PREVAILS AGAINST ODDS
UNITY CONQUERS ALL ADVERSITY AND TRASCENDS ALL MISFORTUNE
UNITY PREVAILS AMIDST ODDS
UNITY CONQUERS ALL ADVERSITY AND TRASCENDS EACH MISFORTUNE
UNITY PREVAILS DESPITE ODDS